OpenStudy (gabylovesyou):

Solve the equation x-6 /2 = 3 for x.

OpenStudy (anonymous):

What if we rewrite -6/2 as -3? Then we just have to add 3 both sides.

OpenStudy (anonymous):

x-6 =2*3 x-6=6 x=??

OpenStudy (anonymous):

x / (x/2-x/6) = x / (3x/6 - x/6) = x / (2x/6) = x * (6/2x) = 6/2 =3

OpenStudy (anonymous):

Is it like this? \[\frac{x-6}{2}=3\] Or this? \[x-\frac{6}{2}=3\]

OpenStudy (anonymous):

yeah

mathslover (mathslover):

which one @Gabylovesyou : please tell

OpenStudy (gabylovesyou):

The first one @rafabc02

OpenStudy (anonymous):

@Gabylovesyou then do as I've done above: x-6 =2*3 x-6=6 x=??

OpenStudy (gabylovesyou):

x = 12??

OpenStudy (anonymous):

x=6

OpenStudy (gabylovesyou):

whyy

OpenStudy (anonymous):